1-(50) Find: RG 15V Mr he - Yas , Im and the region of geration e km = 60 mA/v? V2v IMA 2002 and the region
Answer #1

Solm RD-2001 151 2 RG=lmr m YIGO VG5 Kn=60mA/v2 vi = QV ID NGs ID 2 Is = kn (vers-VT) 2 ID - GOMA /VGs-de -2) 15-200 ID = Vo4.082 - 13-200 a 200 x 2 200x²+4.08 x-13 = 0 4.08 +504.88924200) (-13) 2= 2x 200 2 = -4.087./16.646 + 10400 yo 2- -4.08 + 102gm = Goma gm = kn (NGs-V7) Goma (3-2) Ane [ Im = 60 mAh